OpenStudy (anonymous):

x² – 81x = 19 1; 19 –1; 19 3; 6 –3; 1

OpenStudy (anonymous):

None of the solutions offered are correct. x² – 81x = 19 x² – 81x -19 = 0 x = (81 + √6637)/2, (81 - √6637)/2 x = 81.234, -.234

OpenStudy (anonymous):

Confirmed, none of the solutions are correct.

OpenStudy (anonymous):

question is incorrect

OpenStudy (anonymous):

Or your teacher is a terrible person, and likes tricking students.

OpenStudy (anonymous):

it says on my paper "If necessary, round to the nearest hundredth." does that help any

OpenStudy (anonymous):

No; the closest integer values are \(x=0,80\).

OpenStudy (anonymous):

to your previous answer, i do agree my teacher flutters up alot

OpenStudy (anonymous):

lol it mestted it up F*cks up is what i mean't to say, but thanks i will tall to my teacher